OpenStudy (anonymous):

simplify sec^2A (1-sin^2A)

OpenStudy (anonymous):

what did you do so far?

OpenStudy (anonymous):

im thinking of using the Pythagorean identity 1+tan^2u=sec^2u

OpenStudy (anonymous):

typically, you want to simplify things... going that way gives you more terms. see if there is something to take more terms into less terms

OpenStudy (anonymous):

im stuck :(

OpenStudy (anonymous):

look at the other thing... can you make that into something else?

OpenStudy (anonymous):

sec^2x=1+tan^2x

OpenStudy (anonymous):

no, look at the other term... (1-sin^2A)

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

here's another cheat sheet

OpenStudy (anonymous):

\[1-\sin^2\left( A \right)=?\]\[\sec \left( A \right)=?\]

OpenStudy (anonymous):

\[\sec \left( A \right)=\frac{ 1 }{ \cos \left( A \right) } \Rightarrow \sec \left( A \right)^2=\frac{ 1 }{ \cos \left( A \right)^2 }\]

OpenStudy (anonymous):

oops, i meant\[\sec^2\left( A \right)=\frac{ 1 }{\cos^2\left( A \right) }\]

OpenStudy (anonymous):

also,\[1-\sin^2\left( A \right)=\cos^2\left( A \right)\]

OpenStudy (anonymous):

you need to know and understand these fundamental relationships... The reiprocal functions of cosine, sine and tangent:\[\sec \left( x \right)=\frac{ 1 }{ \cos \left( x \right) }\]\[\csc \left( x \right)=\frac{ 1 }{ \sin \left( x \right) }\]\[\cot \left( x \right)=\frac{ \cos \left( x \right) }{\sin \left( x \right) }\] the Pythagorean Identities: \[\cos^2\left( x \right)+\sin^2\left( x \right)=1\] \[\tan^2\left( x \right)+1=\sec^2\left( x \right)\] \[\cot^2\left( x \right)+1=\csc^2\left( x \right)\]

OpenStudy (anonymous):

I got one as the answer

OpenStudy (anonymous):

that b correctamundo

OpenStudy (anonymous):

yayy (: